OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 4x – 2y = 7 3x + y = 5

OpenStudy (anonymous):

Given  we find (Dx) with 4(1) - (3)(-2) = 4 + 6 = 10

OpenStudy (anonymous):

Thank you so much!! <3